OpenStudy (anonymous):

8 – 4(x – y) = 2x + 6 How to solve this? Please don't give me the answer, want to know how to solve it.

OpenStudy (across):

Begin by distributing the term -4(x-y).

OpenStudy (anonymous):

8 - 4(x - y) = 2x + 6 8 - 4x - 4y = 2x + 6

OpenStudy (anonymous):

like that?

OpenStudy (across):

Almost!

OpenStudy (across):

Look at your second term in that distribution: -4y. What is (-4)*(-y)?

OpenStudy (anonymous):

ohhh, so it'd be... +(positive) 4y?

OpenStudy (across):

Correct! So we now have 8 - 4x + 4y = 2x + 6. The next step is to combine like terms. Do you know how to do that?

OpenStudy (anonymous):

I think so: 8 - 4x - 2x + 4y = 6?

OpenStudy (across):

Well done! Do you agree that -4x-2x=-6x?

OpenStudy (anonymous):

Yes, but how did we get to -6x? What about the 8? (Math is not my best subject)

OpenStudy (across):

We're just simplifying right now: 8 - 4x - 2x + 4y = 6 8 - 6x + 4y = 6

OpenStudy (across):

By the way, are we trying to solve for y?

OpenStudy (anonymous):

oohh, I thought that when you put the equation, that that was the final equation. I'm supposed to just write the equation in the simplest form, but I want to also know how to solve it, so I have back up in case I get a question like that.

OpenStudy (across):

Okay. :) Now that we have 8 - 6x + 4y = 6, our next step would be to combine the remaining two "like" terms 8 and 6. Could you do that?

OpenStudy (anonymous):

8 - 6x + 4y - 8 = 6 - 8 So, it would end up as, -6x + 4y = -2 correct?

OpenStudy (across):

That's right. :) Now, the expression can't be simplified any further, but you could rewrite it as much as you want. For example, you could divide both sides by two to obtain -3x+2y=-1.

OpenStudy (anonymous):

Ohh, my answers are A. 6x – 4y = –2 B. 6x – 4y = 2 C. 4x – 6y = 2 D. 6x + 4y = 2 how do I turn the six into positive?

OpenStudy (across):

Ah, I see. :) Well, we ended up with -6x + 4y = -2. If we multiply both sides by -1, what would we end up with?

OpenStudy (anonymous):

6x + 4y x -1 = -2 x -1 so that would be, 6x + 4y = 2!

OpenStudy (across):

Very close!

OpenStudy (anonymous):

Or do I have to multiply the 4y too?

OpenStudy (across):

What happens to the + 4y term?

OpenStudy (across):

Yep!

OpenStudy (anonymous):

that means it's 6x - 4y = 2?

OpenStudy (across):

There you go. :)

OpenStudy (anonymous):

Yes! Thanks a bunch, math was never my favorite subjects :P Thanks again

OpenStudy (across):

You're most welcome!